OpenStudy (anonymous):

Solve: w(6w + 12) = 0

OpenStudy (dream95):

w=0 or w=-2

OpenStudy (ajprincess):

w=0/6w=-12=-12/6=-2

OpenStudy (callisto):

For this equation to exist, either w=0 or (6w + 12) = 0 So, solve them. and you'll get w=0 or 6w+12=0 6w =-12 w=-2 The solutions to the question are w=0 or w=-2

OpenStudy (anonymous):

w=0 6w+12=0 6w= -12 w = -2

OpenStudy (callisto):

@tana I think you should learn how to do this type of questions instead of just asking so many similar questions